OpenStudy (anonymous):

find the domain of the function. g(x)=1-2x^2

OpenStudy (anonymous):

\[(-\infty,\infty)\]

OpenStudy (anonymous):

There are no limitations on g(x)

OpenStudy (anonymous):

which means all values of x make sense

OpenStudy (anonymous):

I am with Largange's Son

OpenStudy (anonymous):

how about g(y)=radical y-10 ?

OpenStudy (anonymous):

you have to set this up: \[y-10\ge0\]

OpenStudy (anonymous):

now solve like a normal equation: y-10=0, add 10 to both sides and you get y=10 or in other words, y had to be greater than 10

OpenStudy (anonymous):

\[[10,\infty)\]